Suppose the mean expenditure per customer at a tire store is $80.00, with a standard deviation


of $10.00. If a random sample of 40 customers is taken, what is the probability


a) that the sample average expenditure per customer for this sample will be $87.00 or


less?


b) that the sample average expenditure per customer for this sample will be $87.00 or


more?


c) that the sample average expenditure per customer for this sample will be between


$70.00 and $85.00?



1
Expert's answer
2021-07-06T12:13:00-0400

"\mu = 80.00 \\

\sigma=10.00 \\

n=40"

a)

"P(\bar{x} ≤ 87.00) = P(Z≤ \frac{87.00-80.00}{10.00/ \sqrt{40}}) \\

= P(Z≤ 4.42) \\

= 0.9999"

b)

"P(\bar{x} ≥87.00) = 1 -P(Z<87.00) \\

= 1 - 0.9999 \\

= 0.0001"

c)

"P(70.00< \bar{x}<85.00) = P(\bar{x}<85.00) -P(\bar{x}<70.00) \\

=P(Z< \frac{85.00 -80.00}{10.00/ \sqrt{40}}) -P(Z< \frac{70.00 -80.00}{10.00/ \sqrt{40}}) \\

=P(Z<3.16) -P(Z< -6.32) \\

= 0.99921 -0.00003 \\

= 0.99918"